# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v3

@@ 5.2. Adjudicator
Social welfare functions also allow for assigning a different importance to clauses: The user could submit an importance score $w_{i}$ for each clause $p_{i}$, which can be easily incorporated in the social welfare function, e.g., the Utilitarian welfare function becomes $\operatorname*{arg\,max}_{R\in\mathcal{R}}\sum_{i\in[\ell]}w_{i}\cdot s_{i}(R)$.
Selecting the social welfare function gives us control over the tradeoffs between objectives: By picking the egalitarian function, we ensure that one clause will not get prioritized over another. In contrast, the Utilitarian function prioritizes the summed alignment, allowing for mismatches between clauses; the Nash function strikes a balance between the two functions. Further, the adjudicator makes the selection process more transparent, as the different objectives, the selection criterion, and the performance of the candidate reward functions regarding the objectives become explicit.
@@
For each preference clause $p_{i}\in P$, we compute a reward function $R_{i}$ aligned with $p_{i}$ by casting it as a singular prompt to the DLM pipeline (see Section B.2). For each $R\in\mathcal{R}$, we compute as $s_{i}(R)$ the expected reward according to reward function $R_{i}$ produced by policy $\Pi(R)$ (again, we approximate this quantity by running multiple simulations). $s_{i}(R)$ quantifies the quality of the policy induced by the candidate reward function $R$ from the perspective of $p_{i}$ (as captured by $R_{i}$). As the scale of the reward functions can vary significantly among preference clauses, we normalize the scores by the performance of the default policy, i.e., we compute $\frac{s_{i}(R)-s_{i}(R^{*})}{s_{i}(R^{*})}$.
@@
The Simulator Scorer Model assumes access to reward functions capturing individual preference clauses well. If no well-aligned reward functions can be obtained, the performance of SCLM-SIM can deteriorate because it can become very noisy. Another disadvantage of SCLM-SIM is that the scores in SCLM-SIM are all computed via simulation, which can become computationally quite costly. Motivated by this, we propose a quicker and more flexible LLM-based approach, where we prompt an LLM to rate the alignment of a candidate reward function with a preference clause. In particular, for each $R\in\mathcal{R}$ and $p_{i}\in P$, we use a prompt that includes $R$, $p_{i}$, and the utility feature distributions produced by policy $\Pi(R)$. We ask the LLM to rank how well $R$ aligns with the preference clause $p_{i}$ on a scale from 1 to 5 (see Appendix D for prompt texts).
@@
Aligning reward functions to a prioritization prompt may cause (unintended) utility shifts in other features (e.g., due to feature correlations, shifting utility to low-income beneficiaries might shift it away from more educated ones). See Figure 2 for a concrete example. SCLM offers users the option to explicitly prevent these shifts by adding additional clauses (“objectives”) to the prompt: Given a prompt $P$ (e.g., the prompt from Section 4), for each feature not referenced in the prompt, the user can add a new preference clause requesting a minimum shift in the utility distribution of this feature (e.g., for Section 4 they could add “do not change the utility distribution for feature $C$”). To compute the alignment score $s_{i}(R)$ between a reward function $R$ and a clause $p_{i}$=“minimize utility shift for feature $X$”, we compare feature $X$’s utility distribution under the default policy with its utility distribution under the policy $\Pi(R)$. Specifically, we quantify the difference using the Earth mover’s distance (EMD) between the two distributions. Afterward, we apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$, which are input to the social welfare function (along with the alignment scores for the other clauses).
@@
Another potential risk of aligning a reward function with a prioritization prompt is that it can sharply decrease the summed utility generated by all agents: The user might request the prioritization of a subpopulation that does not benefit much from receiving a resource, leading to severe drops in the summed utility generated by all agents. Users can address this issue in our model by adding a clause $p_{i}$=“maximize the total generated utility” to the prompt. As the alignment score $s_{i}(R)$ of $p_{i}$ with some reward function $R$ we compute the summed utility generated by all agents under the policy $\Pi(R)$ (computed via multiple simulations of the policy on the given instance). We again apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$.
6. Experiments
@@
We describe our testset (Section 6.1), the compared methods (Section 6.2), and our experimental results both for dealing with composite prioritization prompts (Section 7.1) and additionally minimizing unintended side effects (Section 5.2). Following the work of Behari et al. (2024), which constitutes our most important baseline, we use Gemini Pro Anil et al. (2023) as the LLM in our experiments.
6.1. Dataset Description
Synthetic Domain
We create three synthetic datasets, each with three features ($A$, $B$, and $C$). For some agent and feature, we randomly sample the agent’s feature value between 00 and $1$. Arm’s passive transition probabilities ($P(s,a=0,s^{\prime})$ for $s,s^{\prime}\in[0,1]$) are sampled uniformly between 00 and $1$. Our three datasets differ in how active transition probabilities are sampled. For each dataset, we set a three-dimensional weight vector $W\in[0,1]^{3}$ specifying how much each feature impacts the effect of applying an active action (see Section B.3 for details). For each agent, let $f$ denote their feature values; we sample their active transition probabilities for $s,s^{\prime}\in[0,1]$ as $P(s,a=1,s^{\prime})=P(s,a=0,s^{\prime})+\delta,\;\delta\sim\mathcal{N}(\Delta,\sigma),$ where $\Delta=W\cdot f^{T}$ and $\sigma$ is a hyperparameter. Subsequently, we discretize the values of all features into $5$ different equal-sized buckets to be consistent with the real-world domain. For each dataset, we sample $5$ instances according to the above procedure with $N=2100$ arms, a budget of $B=210$, and a time horizon of $T=12$ (to replicate the setup of the real-world domain described below where roughly $3\%-10\%$ of beneficiaries receive a resource every week for up to three months).
Real-World Domain
@@ 6.1. Dataset Description
ARMMAN (2024) is a non-profit in India that operates large-scale Maternal and Child Care Mobile Health programs for underserved communities. One of their programs disseminates critical health information via weekly automated voice messages. The goal of the NGO is to maximize beneficiaries’ engagement, i.e., the number of messages they listen to. A limited number of beneficiaries are called by health workers every week to boost engagement. The problem of planning which beneficiaries to call has been modeled and solved as an RMAB, where the good/bad state corresponds to a high/low weekly engagement of the beneficiary. We use anonymized data from a quality improvement study conducted in January 2022 Verma et al. (2023). For each beneficiary, we have access to their income, education, and age level, which we use as our three features. Beneficiaries’ historic listenership values are used to estimate their transition probabilities under the passive action Mate et al. (2022). One problem in estimating transition probabilities under the active action is that due to the limited number of service calls made, such transitions are rare. Thus, active transition probability estimates are noisy. To alleviate this issue, we use the features and passive transition probabilities from ARMMAN together with the procedure to sample active transitions described in the synthetic domain. Again, we create datasets for three different weight vectors $W$. Each dataset consists of five sampled RMAB instances with $N=2100$ arms, a budget of $B=210$ and a time horizon of $T=12$.
@@
6.2. Models & Baselines
We analyze six different variants of SCLM differing in the used social welfare function (Utilitarian, Egalitarian, Nash) and scorer model (Simulator or LLM), e.g., we denote as SCLM-SIM-Egalitarian SCLM with the Simulator Scorer Model and the Egalitarian social welfare function. In our generator, we generate $4$ candidate reward functions in each step and run $5$ iterations to generate a total of $20$ candidate reward functions. In addition, we consider several LLM-focused baselines (see Appendix D for detailed descriptions):
LLM-Zeroshot:
@@ 6.2. Models & Baselines
DLM:
This baseline implements the Decision-Language Model by Behari et al. (2024) (see Section B.2).
DLM-PromptEngg:
This is a modified version of DLM where within the reflection prompt, we include examples for singular queries of how the LLM should reason over the different reward function choices (see Appendix D).
7. Results
7.1. Composite Prioritization Prompts
We analyze the performance on the $12$ composite prompts which request the prioritization of two subpopulations (see Section C.1 for additional results).
@@ 7.1. Composite Prioritization Prompts
Figure 4. Results comparing the quality of reward design methods for composite prioritization prompts. Results are averaged across $180=12\cdot 15$ values: $12$ composite prompts on $15$ RMAB instances (from $3$ datasets). Error bars represent std-error.
@@ Results
If we replace the Simulation Scorer with the LLM Scorer, the performance of the SCLM decreases, but is still better than all of our three baselines. The difference between the LLM and Simulation Scorer highlights the advantage of the additional information acquired through the more complex and computationally expensive simulation method. Regarding the performance of the baselines, our DLM baseline with prompt engineering DLM-PromptEngg improves slightly upon the results of DLM in the synthetic domain, while in the real-world domain, their performance is similar. This suggests that prompt engineering itself is not sufficient to adequately deal with the multi-objective nature of composite prompts; an external component (like our adjudicator) is needed. Finally, LLM-zeroshot consistently performs the worst, which highlights the non-triviality of the reward design problem and the need for a guided extensive search within the reward function search space.
Effect of Social Welfare Function
Figure 5. Comparison of different methods using the prompt “Prioritize agents with a low value for feature $B$ and agents with a low value for feature $C$”.
While both the Utilitarian and Egalitarian social welfare function leads to good overall results, the choice between them significantly influences which reward function is selected on the instance level. Figure 5 shows one example instance for a prompt with two prioritization clauses. As in Figure 1, each point corresponds to a candidate reward function with the axes quantifying the alignment with the two respective clauses (as described in the beginning of this section). The SCLM-Utilitarian model chooses a reward function from the Pareto frontier that shows a much stronger effect for the second clause (i.e., the utility increase for agents with a low value of feature $C$ is more pronounced). In contrast, the reward function selected by SCLM-Egalitarian is much more balanced in distributing the utility increase between the two groups. DLM-based baselines fail to pick a reward function from the Pareto frontier.
Table 1. Results comparing different reward function selection strategies, aggregated across the three real-world datasets. A higher summed % change in desired feature(s) implies a better alignment with the prioritization clauses, whereas less unintended shift and less % drop in utility are better.
@@
(a) Tradeoff between prioritization prompts alignment and unintended shift minimizationy
Method
Summed % Change in
Desired Feature(s)
Unintended Shift
% Drop in
Utility
DLM-PrioritizationOnly 6.809$\pm$0.86 0.302$\pm$0.02 6.233$\pm$0.17
DLM-ExtendedPrompt-Util 3.416$\pm$1.04 0.241$\pm$0.02 5.689$\pm$0.25
SCLM-PrioritizationOnly 13.131$\pm$0.86 0.316$\pm$0.02 5.372$\pm$0.19
SCLM-ExtendedPrompt-Util 8.869$\pm$0.73 0.323$\pm$0.02 3.91$\pm$0.21
SCLM-Full 11.496$\pm$0.87 0.142$\pm$0.01 3.85$\pm$0.18
(b) Tradeoff between prioritization prompts alignment and total utility maximization
7.2. Addressing Fairness and Biases
As discussed in Section 5.2, we can also use our pipeline to prevent unintended side-effects of aligning reward functions with prioritization clauses: In this section, we will focus on preventing shifts in utility feature distribution of features not included in the prompt.
We analyze all $6$ singular and $12$ composite prioritization prompts (see Section 6.1), where we add additional clauses to prevent shifts in the utility distribution of all features not referenced in the prompt. We use the simulator scorer with a utilitarian social welfare function and call the resulting model SCLM-ExtendedPrompt-Fair. As baselines, we consider DLM only prompted with the prioritization clause(s) (called DLM-PrioritizationOnly) and DLM prompted with the prioritization clause(s) and clause(s) asking for the minimizing of utility shifts for the other features (called DLM-ExtendedPrompt-Fair). We also consider SCLM-SIM-Utilaterian only prompted with the prioritization clause(s) (called SCLM-PrioritizationOnly). See Appendix D for more details of prompts.
@@ 7.2. Addressing Fairness and Biases
Figure 5(a) shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Fair, we find that adding additional objective to the prompt does not result in a better performance for real-world domains. In contrast, SCLM-ExtendedPrompt-Fair which incorporates unintended shifts in the selection chooses reward functions resulting in significantly higher utility increases for prioritized subpopulations and significantly fewer unintended utility shifts. The fact that SCLM performs advantageously for both (conflicting) objectives highlights the quality of the pipeline and its capabilities to effectively address multiple objectives (of different types). We see similar results in synthetic domain (see Table 5 in the appendix).
7.3. Preventing Ineffective Resource Use
Figure 5(b) shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Util, we again observe that adding the additional objective to the prompt does not result in a better performance in real-world domains (for similar results on the synthetic domain, see Table 6 in the appendix). In contrast, SCLM-ExtendedPrompt-Util which incorporates utility as a scoring dimension chooses reward functions resulting in a significantly higher utility increases for the prioritized subpopulations and a significantly higher total utility (i.e., there is a lower drop in utility).
Combining Prioritization, Unintended Shifts, and Utility Drops
Finally, we consider a variant of SCLM which both prevents unintended shifts and utility drops, i.e., in addition to the specified prioritization clauses, we add clause(s) on the minimization of shifts and a clause on the minimization of utility drops (which results in having a total of 4 evaluation dimensions; one per feature plus the utility drop one). We use the simulator scorer together with the utilitarian welfare function and call the resulting model SCLM-Full. We see in the last row of Figure 5(b) that this model chooses balanced reward functions with high % change in utility in desired attributes, high overall utility, and low unintended shift, showcasing that SCLM can flexibly allow the end-user to combine multiple evaluation dimensions.
8. Discussion
We present a customizable Social Choice Language Model to handle the multi-objective nature of preference prompts in reward design for RMABs. We showcase how methods from social choice can be used to improve the quality and transparency of decision-making of LLM-based frameworks, as we present an adjudicator component that makes the final decision from options generated by the LLM. SCLM significantly improves the quality of the chosen reward functions. We demonstrate that SCLM can not only handle composite prioritization prompts but arbitrary prompts containing multiple objectives, e.g., balancing the prioritization of subpopulations with the total utility generated by all agents. For future work, SCLM can be applied to other problems from multiagent planning and reinforcement learning. Further, SCLM can easily be extended to handle multiple preference prompts specified by different users. As our model is compatible with any social welfare function, it would also be possible to let the user manually configure the social welfare function. One could in particular allow for adding weights in front of the different alignment scores, enabling the user to express the importance of different clauses (respectively, objectives).

### v4

@@ 5.2.1 Selection via Social Welfare Function
Selecting the social welfare function gives us control over the tradeoffs between objectives: By picking the Egalitarian function, we ensure that one clause will not get prioritized over another. In contrast, the Utilitarian function prioritizes the summed alignment, allowing for mismatches between clauses; the Nash function strikes a balance between the two functions. [^1] Note that social welfare functions also allow for assigning a different importance to clauses: The user could submit an importance score $w_{i}$ for each clause $p_{i}$, which can be easily incorporated in the social welfare function, e.g., the Utilitarian welfare function becomes $\operatorname*{arg\,max}_{R\in\mathcal{R}}\sum_{i\in[\ell]}w_{i}\cdot s_{i}(R)$. The adjudicator makes the selection process more transparent, as the different objectives, the selection criterion, and the performance of the candidate reward functions regarding the objectives become explicit.
5.2.2 Computing Alignment Scores
@@
For each preference clause $p_{i}\in P$, we compute a reward function $R_{i}$ aligned with $p_{i}$ by casting it as a singular prompt to the DLM pipeline (see Section C.2). For each $R\in\mathcal{R}$, we compute as $s_{i}(R)$ the expected reward according to reward function $R_{i}$ produced by policy $\Pi(R)$ (again, we approximate this quantity by running multiple simulations). Accordingly, $s_{i}(R)$ quantifies the quality of the policy induced by the candidate reward function $R$ from the perspective of $p_{i}$ (as captured by $R_{i}$). As the scale of the reward functions can vary significantly among preference clauses, we normalize the scores by the performance of the default policy, i.e., we compute $\frac{s_{i}(R)-s_{i}(R^{*})}{s_{i}(R^{*})}$.
@@
The Simulator Scorer Model assumes access to reward functions capturing individual preference clauses well. If no well-aligned reward functions can be obtained, the performance of SCLM-SIM can deteriorate because it can become noisy. Another disadvantage of SCLM-SIM is that the scores in SCLM-SIM are all computed via simulation, which can become computationally costly. Motivated by this, we propose a quicker and more flexible LLM-based approach, where we prompt an LLM to rate the alignment of a candidate reward function with a preference clause. In particular, for each $R\in\mathcal{R}$ and $p_{i}\in P$, we use a prompt that includes $R$, $p_{i}$, and the utility feature distributions produced by policy $\Pi(R)$. We ask the LLM to rank how well $R$ aligns with the preference clause $p_{i}$ on a scale from 1 to 5 (see Appendix F for prompt texts).
@@
Aligning reward functions to a prioritization prompt may cause (unintended) utility shifts in other features (e.g., due to feature correlations, shifting utility to low-income beneficiaries might shift it away from more educated ones). See Figure 5 for a concrete example. SCLM offers users the option to explicitly prevent these shifts by adding additional clauses (“objectives”) to the prompt: Given a prompt $P$ (e.g., the prompt from Example 1), for each feature not referenced in the prompt, the user can add a new preference clause requesting a minimum shift in the utility distribution of this feature (e.g., for Example 1 they could add “do not change the utility distribution for feature $C$”). To compute the alignment score $s_{i}(R)$ between a reward function $R$ and a clause $p_{i}$=“minimize utility shift for feature $X$”, we compare feature $X$’s utility distribution under the default policy with its utility distribution under the policy $\Pi(R)$. Specifically, we quantify the difference using the Earth mover’s distance (EMD) between the two distributions. Afterward, we apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$, which are input to the social welfare function (along with the alignment scores for the other clauses).
@@
Another potential risk of aligning a reward function with a prioritization prompt is that it can sharply decrease the summed utility generated by all agents: The user might request the prioritization of a subpopulation that does not benefit much from receiving a resource, leading to severe drops in the summed utility generated by all agents. Users can address this issue in our model by adding a clause $p_{i}$=“maximize the total generated utility” to the prompt. As the alignment score $s_{i}(R)$ of $p_{i}$ with some reward function $R$ we compute the summed utility, i.e., the total number of steps in which arms are in an active state, generated by all agents under the policy $\Pi(R)$ (computed via multiple simulations of the policy on the given instance). We again apply 00-$1$ normalization to all scores $s_{i}(R)_{R\in\mathcal{R}}$ for prompt $p_{i}$.
5.2.3 Error Bounds for Adjudicator’s Selection
Even though we observe in our experiments that the scorer models produce mostly accurate scores, the output scores are oftentimes still a bit noisy. To measure how errors propagate through the Social Choice Language Model and how they affect the final reward function selection, we consider the following setup.
Suppose instead of observing the true score vector $\mathbf{s}(R_{j})=(s_{1}(R_{j}),s_{2}(R_{j}),...,s_{l}(R_{j}))$, the Scorer Model (SCLM-SIM or SCLM-LLM) returns a noisy score estimate $\tilde{\mathbf{s}}(R_{j})$ with multiplicative noise $\alpha\in(0,1]$ satisfying
$\alpha\cdot\mathbf{s}(R_{j})\leq\tilde{\mathbf{s}}(R_{j})\leq\frac{1}{\alpha}\cdot\mathbf{s}(R_{j}),\quad\forall R_{j}\in\mathcal{R}.$ (2)
Let $\tilde{R}^{*}=\arg\max_{R_{j}\in\mathcal{R}}f_{p}(\tilde{\mathbf{s}}(R_{j}))$ be the best reward function under the generalized p-mean function for the observed, noisy scores ($\tilde{R}^{*}$ will be returned by SCLM); and $R^{*}=\arg\max_{R_{j}\in\mathcal{R}}f_{p}(\mathbf{s}(R_{j}))$ be the best reward function under the generalized p-mean function for the true, latent score. We define the (relative) regret we encounter by choosing $\tilde{R}^{*}$ instead of $R^{*}$ as:
Relative $\textit{Relative Regret}=\frac{f_{p}(\mathbf{s}(R^{*}))-f_{p}(\mathbf{s}(\tilde{R}^{*}))}{f_{p}(\mathbf{s}(R^{*}))}$ (3)
The relative regret measures the relative drop in p-mean welfare of the reward function chosen by the adjudicator as compared to the optimal reward function. We show that the relative regret degrades gracefully in the multiplicative error parameter, highlighting that even in the presence of noise, SCLM selects good reward functions with guarantees.
Proposition 1.
The relative regret is bounded by $1-\alpha^{2}$.
Proof sketch. We observe the monotonicity and positive homogeneity of the generalized p-mean function applied to the input values. These properties allow the application of function $f$ to both sides of Inequality 2. Subsequently, using the definition of $R^{*}$ and $\tilde{R}^{*}$, we derive the regret bound. For a complete proof, see Appendix D.
6 Experiments
@@
We describe our testset (Section 6.1), the compared methods (Section 6.2), and our experimental results both for dealing with composite prioritization prompts (Section 6.3) and additionally minimizing unintended side effects (Section 6.4). Following the work of Behari et al. [7], which constitutes our most important baseline, we use Gemini Pro [2] as the LLM in our experiments.
6.1 Dataset Description
@@
ARMMAN [3] is a non-profit in India that operates large-scale Maternal and Child Care Mobile Health programs for underserved communities. One of their programs disseminates critical health information via weekly automated voice messages. The goal of the NGO is to maximize beneficiaries’ engagement, i.e., the number of messages they listen to. A limited number of beneficiaries are called by health workers every week to boost engagement. The problem of planning which beneficiaries to call has been modeled and solved as an RMAB, where the good/bad state corresponds to a high/low weekly engagement of the beneficiary. We use anonymized data from a quality improvement study conducted in January 2022 [32]. For each beneficiary, we have access to their income, education, and age level, which we use as our three features. Beneficiaries’ historic listenership values are used to estimate their transition probabilities under the passive action [18]. One problem in estimating transition probabilities under the active action is that due to the limited number of service calls made, such transitions are rare. Thus, active transition probability estimates are noisy. To alleviate this issue, we use the features and passive transition probabilities from ARMMAN together with synthetically generated active transition probabilities. Finally, we create three datasets, each consisting of five sampled RMAB instances with $N=2100$ arms, a budget of $B=210$ and a time horizon of $T=12$. The three datasets differ in how much each feature impacts the effect of applying an active action. In addition to the real-world domain, we also create three completely synthetic domain datasets (see Appendix B.3 and B.4 for more details on dataset generation).
@@
6.2 Models & Baselines
We analyze six different variants of SCLM differing in the used social welfare function (Utilitarian, Egalitarian, Nash) and scorer model (Simulator or LLM), e.g., we denote as SCLM-SIM-Egalitarian SCLM with the Simulator Scorer Model and the Egalitarian social welfare function. In our generator, we generate $4$ candidate reward functions in each step and run $5$ iterations to generate a total of $20$ candidate reward functions. In addition, we consider several LLM-focused baselines (see Appendix F for detailed descriptions):
LLM-Zeroshot
@@ 6.2 Models & Baselines
DLM
This baseline implements the Decision-Language Model by Behari et al. [7] (see Section C.2).
DLM-PromptEngg
This is a modified version of DLM where within the reflection prompt, we include examples for singular queries of how the LLM should reason over the different reward function choices (see Appendix F).
6.3 Results: Composite Prioritization Prompts
We analyze the performance on the $12$ composite prompts described above which request the prioritization of two subpopulations (see Section E.1 for additional results).
@@ 6.3 Results: Composite Prioritization Prompts
Figure 4: Results comparing the quality of reward design methods for composite prioritization prompts. Results are averaged across $180=12\cdot 15$ values: $12$ composite prompts on $15$ RMAB instances (from $3$ datasets). Error bars represent std-error.
@@ Results
If we replace Simulation Scorer with LLM Scorer, performance of SCLM decreases, but is better than all of our three baselines. The difference between LLM and Simulation Scorer highlights the advantage of the additional information acquired through more complex and computationally expensive simulation method. Regarding the performance of baselines, our DLM baseline with prompt engineering DLM-PromptEngg improves slightly upon the results of DLM in the synthetic domain, while in real-world domain, their performance is similar. This suggests that prompt engineering itself is not sufficient to adequately deal with the multi-objective nature of composite prompts; an external component (like our adjudicator) is needed. Finally, LLM-zeroshot consistently performs the worst, which highlights the non-triviality of reward design problem and the need for a guided extensive search within reward function search space.
Table 1: Results comparing different reward function selection strategies, aggregated across three real-world datasets. Higher summed % change in desired feature(s) implies better alignment with prioritization clauses, whereas less unintended shift and less % drop in utility are better.
@@
6.4 Addressing Fairness and Biases
As discussed in Section 5.2, we can also use our pipeline to prevent unintended side-effects of aligning reward functions with prioritization clauses, i.e., (i) shifts in the utility feature distribution of features not included in the prompt, (ii) drops in the total generated utility, and (iii) arbitrary weighted combinations of the above two goals and additional preference clauses. We focus on (i) here and relegate the results for (ii) and (iii), which paint a very similar picture, to Appendix D.3.
We analyze all $6$ singular and $12$ composite prioritization prompts (see Section 6.1), where we add additional clauses to prevent shifts in the utility distribution of all features not referenced in the prompt. We use the simulator scorer with a Utilitarian social welfare function and call the resulting model SCLM-ExtendedPrompt-Fair. As baselines, we consider DLM only prompted with the prioritization clause(s) (called DLM-PrioritizationOnly) and DLM prompted with the prioritization clause(s) and clause(s) together for a request for minimizing of utility shifts for the other features (called DLM-ExtendedPrompt-Fair). We also consider SCLM-SIM-Utilaterian only prompted with the prioritization clause(s) (called SCLM-PrioritizationOnly). See Appendix E for more details on the prompts.
@@ 6.4 Addressing Fairness and Biases
Table 1 shows the results. Comparing DLM-PrioritizationOnly and DLM-ExtendedPrompt-Fair, we find that adding additional objective to the prompt does not result in a better performance for real-world domains. In contrast, SCLM-ExtendedPrompt-Fair which incorporates unintended shifts in the selection chooses reward functions resulting in significantly higher utility increases for prioritized subpopulations and significantly fewer unintended utility shifts. The fact that SCLM performs advantageously for both (conflicting) objectives highlights the quality of the pipeline and its capabilities to effectively address multiple objectives (of different types). We see similar results in synthetic domain (see Table 5 in the appendix).
7 Discussion
We present a customizable Social Choice Language Model to handle the multi-objective nature of preference prompts in reward design for RMABs. We showcase how methods from social choice can be used to improve the quality and transparency of decision-making of LLM-based frameworks, as we present an adjudicator component that makes the final decision from options generated by the LLM. SCLM significantly improves the quality of the chosen reward functions. We demonstrate that SCLM can not only handle composite prioritization prompts but arbitrary prompts containing multiple objectives, e.g., balancing the prioritization of subpopulations with the total utility generated by all agents. For future work, SCLM can be applied to other problems from multiagent planning and reinforcement learning. Further, SCLM can easily be extended to handle multiple preference prompts specified by different users.